ROTATIONAL SPECTRA OF THE FREE RADICALS C_{10}H, C_{12}H, C_{13}H, AND C_{14}H IN A SUPERSONIC JET

C. A. GOTTLIEB, \underline{M. C. McCARTHY}, M. J. TRAVERS, AND P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138 and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138.

Four new carbon chain radicals C_{10}H, C_{12}H, C_{13}H, and C_{14}H have been observed in a pulsed supersonic molecular beam with a Fourier transform microwave spectrometer. The radicals were produced in a discharge through a dilute diacetylene/neon mixture in the throat of a supersonic nozzle. All are found to be linear with ^{2}Π electronic ground states, and all except C_{14}H have resolved lambda-type doubling. At least 10 rotational transitions, between 6 and 16 GHz, were measured in the lowest spin component — ^{2}Π_{3/2} of C_{10}H, C_{12}H, and C_{14}H, and the ^{2}Π_{1/2} component of C_{13}H. Only three spectroscopic constants in the standard Hamiltonian for a molecule in a ^{2}Π state were required to reproduce the spectra to a few parts in 10^{7}: an effective rotational constant, a centrifugal distortion constant, and a lambda-type doubling constant. Detection of these highly unsaturated carbon chains establishes that C_{n}H radicals containing up to 14 carbon atoms are readily produced in a supersonic molecular beam. The relative abundance of C_{n}H radicals with an even number of carbon atoms is fairly constant from C_{6}H through C_{12}H. Although the new radicals are about two orders of magnitude less abundant than C_{4}H, the strong predicted ^{2}Π −^{2} Π electronic transitions may be detectable in a supersonic jet by standard laser spectroscopic techniques.